# Leveraging Geometric Priors for Unaligned Scene Change Detection

### v1

@@ Abstract
Unaligned Scene Change Detection aims to detect scene changes between image pairs captured at different times without assuming viewpoint alignment. To handle viewpoint variations, current methods rely solely on 2D visual cues to establish cross-image correspondence to assist change detection. However, large viewpoint changes can alter visual observations, causing appearance-based matching to drift or fail. Additionally, supervision limited to 2D change masks from small-scale SCD datasets restricts the learning of generalizable multi-view knowledge, making it difficult to reliably identify visual overlaps and handle occlusions. This lack of explicit geometric reasoning represents a critical yet overlooked limitation. In this work, we are the first to leverage geometric priors from a Geometric Foundation Model to address the core challenges of unaligned SCD, including reliable identification of visual overlaps, robust correspondence establishment, and explicit occlusion detection. Building on these priors, we propose a training-free framework that integrates them with the powerful representations of a visual foundation model to enable reliable change detection under viewpoint misalignment. Through extensive evaluation on the PSCD, ChangeSim, and PASLCD datasets, we demonstrate that our approach achieves superior and robust performance. Our code will be released at https://github.com/ZilingLiu/GeoSCD.
@@ I INTRODUCTION
Figure 1: Comparison between GFM-based geometric correspondences and optical flow-based correspondences. Invalid correspondence means that the matched point falls outside the image bounds. Leveraging these robust correspondences, our method generates more accurate change masks.
@@
  • We are the first to introduce geometric priors from GFM to fundamentally address the core challenges of unaligned SCD: accurately identifying visual overlaps, establishing robust correspondences and explicitly detecting occlusions, capabilities that are difficult to achieve with 2D visual cues only.
@@ II-B Geometric Foundation Model
Geometric foundation models (GFMs) are attention-based models that directly infer core 3D properties including camera parameters, depth, and per-pixel point maps from images, thereby providing explicit multi-view geometric priors [12]. Recent advances [14, 13, 15] have shifted the focus from robust two-view point-map regression [12, 24] toward scalable multi-view and multi-attribute prediction. Beyond reshaping traditional reconstruction pipelines, GFMs have demonstrated strong impact across diverse downstream tasks. For example, they enable realistic scene and camera control in video generation [25], replace traditional pose optimization with direct point-map regression in SLAM [26], and enhance representation learning for 3D scene understanding [27]. These applications highlight the effectiveness of GFMs’ geometric priors for processing multi-view data and boosting downstream performance. Motivated by these advances, we are the first to adopt GFMs in unaligned SCD to address the key challenges of viewpoint misalignment

### v2

@@ Abstract
Unaligned Scene Change Detection aims to detect scene changes between image pairs captured at different times without assuming viewpoint alignment. To handle viewpoint variations, current methods rely solely on 2D visual cues to establish cross-image correspondence to assist change detection. However, large viewpoint changes can alter visual observations, causing appearance-based matching to drift or fail. Additionally, supervision limited to 2D change masks from small-scale SCD datasets restricts the learning of generalizable multi-view knowledge, making it difficult to reliably identify visual overlaps and handle occlusions. This lack of explicit geometric reasoning represents a critical yet overlooked limitation. In this work, we introduce geometric priors for the first time to address the core challenges of unaligned SCD, for reliable identification of visual overlaps, robust correspondence establishment, and explicit occlusion detection. Building on these priors, we propose a training-free framework that integrates them with the powerful representations of a visual foundation model to enable reliable change detection under viewpoint misalignment. Through extensive evaluation on the PSCD, ChangeSim, and PASLCD datasets, we demonstrate that our approach achieves superior and robust performance. Our code will be released at https://github.com/ZilingLiu/GeoSCD.
@@ I INTRODUCTION
Figure 1: Comparison between GFM-based geometric correspondences and optical flow-based correspondences. Invalid correspondence means that the matched point falls outside the image bounds. Leveraging geometric correspondences, our method generates more accurate change masks.
@@
  • We are the first to introduce geometric priors from GFM to accurately identify visual overlaps, establish robust correspondences, and explicitly detect occlusions, which are difficult to achieve with 2D visual cues only.
@@ II-B Geometric Foundation Model
Geometric foundation models (GFMs) are attention-based models that directly infer core 3D properties including camera parameters, depth, and per-pixel point maps from images, thereby providing explicit multi-view geometric priors [12]. Recent advances [14, 13, 15, 24] have shifted the focus from robust two-view point-map regression [12, 25] toward scalable multi-view and multi-attribute prediction. Beyond reshaping traditional reconstruction pipelines, GFMs have demonstrated strong impact across diverse downstream tasks. For example, they enable realistic scene and camera control in video generation [26], replace traditional pose optimization with direct point-map regression in SLAM [27], and enhance representation learning for 3D scene understanding [28]. These applications highlight the effectiveness of GFMs’ geometric priors for processing multi-view data and boosting downstream performance. Motivated by these advances, we are the first to adopt GFMs in unaligned SCD to address the key challenges of viewpoint misalignment